In a population of lions, the proportionate death rate is 0.55 per year and the

proportionate birth rate is 0.45 per year. Formulate a model of the population. Solve

the model and discuss its long term behavior. Also, find the equilibrium point of the

model.


1
Expert's answer
2021-06-01T17:40:29-0400

Let B be the birth rate=0.45/year

and D be the death rate =0.55/year


Let N be the Population size of Model


Then Rate of propulation Growth


"\dfrac{dN}{dt}=\text {Birt rate - Death rate}=B-D"




"\dfrac{dN}{dt}=0.45-0.55=-0.1/year<0"


Natural growth decay equation:


"\dfrac{dN}{dt}=-0.1N"

Solve the equation


"N(t)=N(0)e^{-0.1t}"

where "N(0)" is the population at time "t=0."


"N(t)\to0" as "t\to \infin."


Since "-0.1N\not=0" for "t\geq0," then there are no equlibrium points.